What is the circumference of a circle with a diameter of 6cm.
Answer by priya(51)

perimeter of the circle is pi* d or 2*pi*r
where r is the radius of the circle,d is the diameter.
the answer is pi*6
= 3.14 * 6 (pi = 22/7 or 3.14)
= 18.84 cm

want maths lessons online..visit my site!!
priya